OpenStudy (megannicole51):

Find the arc length of the curve r=4/(theta) for [pi/2, pi]

OpenStudy (wolf1728):

Is there any diagram with that?

OpenStudy (megannicole51):

nope:/

OpenStudy (wolf1728):

Hmmm so I'm guessing the radius is 4, and theta can be pi/2 and pi.

OpenStudy (megannicole51):

okay...

OpenStudy (wolf1728):

arc length = radius * theta (in radians) arc length = 4 * pi/2 arc length = 6.2831853072

OpenStudy (wolf1728):

arc length = 4* pi arc length = 12.5663706144

OpenStudy (megannicole51):

it says its wrong:/

OpenStudy (wolf1728):

Okay I'll have to find out what pi radians are.

OpenStudy (megannicole51):

okay:)

OpenStudy (wolf1728):

Hmm according to what I found So π radians = 180° Basically 3.14159265 radians = 180°

OpenStudy (nincompoop):

a pi is 180 degrees

OpenStudy (wolf1728):

That question is worded strangely - Find the arc length of the curve r=4/(theta) for [pi/2, pi] so the curve is r / theta?

OpenStudy (nincompoop):

arc length refers to radian @wolf1728 

OpenStudy (megannicole51):

yes